# Homeomorphism groups of basilica, rabbit and airplane Julia sets

### v1

@@ Introduction
Thanks to this uniqueness, we denote by $\mathcal{A}$ the airplane and $\mathcal{R}_{n}$ the $n$-regular rabbit. The key tool to prove A is to associate another topological space that encodes the arrangements of circles: the dendrite or tree of circles. For the airplane, this dendrite is the universal Ważewski dendrite $D_{\infty}$. Let us recall that a dendrite is a Peano continuum where any two points are joined by a unique arc and can be thought as a compact metric tree. The universal Ważewski dendrite is the unique dendrite such that branch points have infinite order and are dense (it is universal in the sense that it includes homeomorphically any other dendrite, see [Wa23]). For the regular $n$-rabbit, the tree of circles is the $(n,\infty)$-biregular tree $T_{n,\infty}$, i.e., any edge joins a vertex of degree $n$ and one of infinite degree.
@@
In [BM00], Burger and Mozes introduced the concept of a universal group $\mathcal{U}(\Gamma)$ for automorphisms of the $n$-regular tree $T_{n}$, given a permutation group $\Gamma\leq\mathrm{Sym}([n])$ (where $[n]=\{1,\dots,n\}$ if $n$ is finite and $[\infty]=\mathbf{N}$). The universal group $\mathcal{U}(\Gamma)$ is characterized by the property that the local action, i.e., the action induced on the neighborhood of each vertex, lies in $\Gamma$. This construction was later generalized by Smith in [Smi17] for biregular trees $T_{m,n}$, where $m,n$ are countable and possibly infinite. In this case, two permutation groups $\Gamma_{m}\leq\mathrm{Sym}([m])$ and $\Gamma_{n}\leq\mathrm{Sym}([n])$ are used to construct the universal group $\mathcal{U}(\Gamma_{m},\Gamma_{n})$.
@@
Figure 5: The $4$-rabbit lamination and the airplane lamination.
@@
The groups $\mathrm{Homeo}^{+}(\mathcal{A})$ and $\mathrm{Homeo}^{+}(\mathcal{R}_{n})$ are the subgroups of $\mathrm{Homeo}^{+}(\mathbf{S}^{1})$ that preserve the associated laminations, which are closed in $\mathrm{Homeo}^{+}(\mathbf{S}^{1})$.
Consequences of the identification of the homeomorphism groups.
@@
The two topological groups $\mathrm{Homeo}(\mathcal{R}_{n})$ and $\mathrm{Homeo}^{+}(\mathcal{R}_{n})$ are locally Roelcke precompact, in particular locally bounded, and quasi-isometric to the infinite-order tree $T_{\infty}$. On the other hand, the topological groups $\mathrm{Homeo}(\mathcal{A})$ and $\mathrm{Homeo}^{+}(\mathcal{A})$ are Roelcke precompact and thus coarsely bounded.
@@
In the case of $\mathrm{Homeo}(\mathcal{A})$, this group can be identified with a kaleidoscopic group where the local action is oligomorphic. As a result, the identification of the universal minimal flow of $\mathrm{Homeo}(\mathcal{A})$ follows directly from [BT23, Theorem 1.1]. For $\mathrm{Homeo}^{+}(\mathcal{A})$, however, the universal minimal flow has a particularly explicit description. Following [GM18a], for a subset $A\subseteq\mathbf{S}^{1}$, let $\mathrm{Split}(\mathbf{S}^{1},A)$ denote the compact space obtained from $\mathbf{S}^{1}$ (with its continuous cyclic order) by replacing each point $a\in A$ with a pair of points $a^{-},a^{+}$, where $a^{+}$ is the cyclic successor of $a^{-}$. Let $P$ be the countable set that corresponds to the preimage of the set of cut points of $\mathcal{A}$ under the Carathéodory loop.
@@
gives an example where the universal minimal flow is cyclically ordered while some other minimal flows are not, since $\mathcal{A}$ and $D_{\infty}$ have no invariant cyclic order under $\mathrm{Homeo}^{+}(\mathbf{S}^{1})$ (Proposition 6.6). This answers a question of Glasner and Megrelishvili, [GM18a, Question 4.9].
@@
The pictures of Julia sets displayed in this article were produced using some Python code developed by the first author. It is available in the following git repository for other uses: https://gitlab.imo.universite-paris-saclay.fr/bruno.duchesne/julia-sets-in-python
@@ Acknowledgements
We thank Matthieu Joseph, Eli Glasner and Michael Megrelishvili and Yuri Neretin for their comments on a preliminary version of this text.
@@ 1.1 Topological characterization of rabbits
Without condition 3, A would include Julia sets where circles accumulate towards certain points which are not end points, such as the Julia set for the map $z\mapsto z^{2}+c$ where $c\approx-1.311$, which is depicted in Figure 6.
@@
  • a regular point if it belongs to a circle and is not a cut point or, equivalently, if it belongs to a unique circle.
@@
Let $I=[a,b]$ be an arc in a rabbit $\mathcal{R}$ and let $p\in(a,b)\coloneqq[a,b]\setminus\{a,b\}$. If $p$ is a cut point, then it belongs to multiple circles by Corollary 1.5, so there is nothing to prove. If $p$ is not a cut point, then $\mathcal{R}\setminus\{p\}$ is connected and thus path-connected because it is a Peano continuum. It follows that there exists an arc $I^{*}$ that joins $a$ and $b$ while avoiding $p$. Then $I\cup I^{*}$ must include some circle that contains $p$. ∎
@@ 1.3 Legal coloring of the tree of circles
Let $\mathcal{R}$ be an $n$-regular rabbit. Let $\Omega$ be a countable set together with a dense cyclic order $O_{\Omega}$ and, if $n$ is finite, let $[n]=\{1,\dots,n\}$ be equipped with its standard cyclic order. For each circle $C$ of $\mathcal{R}$, equip $\mathrm{out}(C)$ with a countable dense cyclic order on the cut points of $C$ as described in Remark 1.19 and, if $n$ is finite, for each cut point $p$ of $\mathcal{R}$, equip $\mathrm{out}(p)$ with a cyclic order. Then there is a legal coloring with colors $([n],\Omega)$ whose restriction to $\mathrm{out}(v)$ is an isomorphism of cyclic orders for each vertex $v$ (only for those that correspond to circles, if $n$ is infinite). In particular, by Lemma A.5 it is an isomorphism of the induced separation relations (Definition A.2).
@@
Suppose that $p$ is a cut point that does not belong to $T$ and is adjacent in $T_{\mathcal{R}}$ to some circle $C$ of $T$. We need to show that we can add $p$ to $T$ along with all circles that are adjacent to it and extend the coloring in such a way that the above conditions are satisfied. This will eventually color every vertex of $T_{\mathcal{R}}$. Given $p$ and $C$ as above, define $k(p,C)$ as $k(q,C)$ for any other cut point $q$ in $T$ that belongs to $C$ (in the first step, when $V(T)=\{C\}$, this choice is instead arbitrary). Then complete $k(p,-)$ to a bijection such that $k(p,C^{\prime})$ satisfies condition 2. Next, for each circle $C^{\prime}\neq C$ that is adjacent to $p$ define $k(C^{\prime},p)=k(C,p)$, so that condition 3 holds at the vertex $p$. For each $C^{\prime}$, color $\mathrm{out}(C^{\prime})$ in such a way that condition 1 holds. The map $k$ extended in this way satisfies all conditions, so we are done. ∎
@@ 1.4 Universal groups of biregular trees
$k|_{\mathrm{out}(g(v))}\circ g|_{\mathrm{out}(v)}\circ k|_{\mathrm{out}(v)}^{-1}(v)\in\begin{cases}N&\text{if $v\in A$},\\ M&\text{if $v\in B$},\end{cases}$ (1)
@@
$\sigma_{g}(v)\coloneqq k|_{\mathrm{out}(g(v))}\circ g|_{\mathrm{out}(v)}\circ k|_{\mathrm{out}(v)}^{-1}(v).$
@@ 1.5 Uniqueness of the regular rabbits
The following proposition shows that there is a unique $n$-regular rabbit for each $n$ (up to homeomorphisms) and is going to be useful later for Proposition 1.24. Given a rabbit $\mathcal{R}$ and its tree of circles $T$, below we denote by $\pi\colon\mathcal{R}\to\overline{T}$ the map that sends each end point to the corresponding point of $\partial T$, each cut point to itself (as a vertex of $T$) and each regular point to the unique circle to which it belongs (as a vertex of $T$).
@@
Let $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ be two $n$-regular rabbits and let $T_{1}$ and $T_{2}$ be their trees of circles. Fix two legal colorings satisfying Lemma 1.21 and let $\varphi\colon T_{1}\to T_{2}$ be a color-preserving automorphism (the local action at any vertex is the identity). Note that the maps $\pi_{1}$ and $\pi_{2}$ are injective on the sets of cut points of $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$. This defines a bijection $\psi=\pi^{-1}_{2}\circ\varphi\circ\pi_{1}$ between the cut points of $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$. Since $\varphi$ is a tree isomorphism and the edge adjacency in $T_{1}$ and $T_{2}$ is given by the inclusion of cut points to circles, the map $\psi$ restricts to bijections between the cut points of each circle. Moreover, since $\varphi$ is color-preserving, the circular order of cut points on each circle is preserved, and since the sets of cut points are dense in $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ (by condition 1 of A together with Lemma 1.21), $\psi$ extends uniquely to a continuous map between a circle of $\mathcal{R}_{1}$ and its image. Doing this for all circles of $\mathcal{R}_{1}$ extends $\psi$ to a bijection between $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ such that $\psi\circ\pi_{1}=\pi_{2}\circ\psi$. It now remains to show that $\psi$ is a homeomorphism.
@@ 1.6 Homeomorphism groups of regular rabbits
Since both are Polish groups, it suffices to prove that $\Pi\colon\mathrm{Homeo}(\mathcal{R})\to\mathcal{U}(\mathrm{Sym}([n]),\mathrm{Aut}(S))$ is continuous (for example because of [Gao09, Corollary 2.3.4]). Suppose that $g_{m}$ is a sequence in $\mathrm{Homeo}(\mathcal{R})$ that converges to the identity and consider an arbitrary circle $C_{0}$ of the rabbit $\mathcal{R}$. We will show that, for large enough $m\in\mathbf{N}$, the homeomorphism $g_{m}$ fixes $C_{0}$. Let $C_{1}$, $C_{2}$ and $C_{3}$ be circles of $\mathcal{R}$ that lie in three distinct connected components $U_{1}$, $U_{2}$ and $U_{3}$ of $\mathcal{R}\setminus C_{0}$. Then, for $m$ large enough, say $m\geq N$, the circle $g_{m}(C_{i})$ is included in $U_{i}$ for each $i=1,2,3$. In the tree of circles, consider the unique paths between $C_{1}$ and $C_{2}$, between $C_{1}$ and $C_{3}$ and between $C_{2}$ and $C_{3}$: by Remark 1.13, they intersect precisely at $C_{0}$, so the vertex $C_{0}$ is fixed by $\Pi(g_{m})$ for all $m\geq N$. This argument shows that each vertex is fixed by $\Pi(g_{m})$ for large enough $m$, so the sequence $\Pi(g_{m})$ converges to the identity of $\mathcal{U}(\mathrm{Sym}([n]),\mathrm{Aut}(S))$ and thus $\Pi$ is continuous. ∎
@@ 1.7 Consequences of the identification with universal groups
Consider now the case $n=2$, which concerns the basilica $\mathcal{R}_{2}=\mathcal{B}$. Neretin noted that $\mathrm{Homeo}(\mathcal{B})$ is the universal Burger-Mozes group $\mathcal{U}(\mathrm{Aut}(S))$ on a regular tree [Ner24], which has an index-$2$ simple subgroup generated by the edge stabilizers by [BM00, Proposition 3.2.1]. The regular tree in question is a simplified tree of circles whose vertices are the circles of the basilica and whose edges connect adjacent circles. Hence, the group generated by the pointwise stabilizers of edges corresponds to the group that preserves the alternating $2$-coloring of the circles depicted in Figure 8. This is the kernel of the group homomorphism $f\colon\mathrm{Homeo}(\mathcal{B})\to\mathrm{Sym}([2])=\{0,1\}$ that has value 00 or $1$ depending on whether the element preserves the coloring or not, which implies that $\mathrm{Homeo}(\mathcal{B})^{\prime}\leq\mathrm{Ker}(f)$. Moreover, note that $\mathrm{Ker}(f)=\mathcal{U}(1,\mathrm{Aut}(S))$, so $\mathrm{Ker}(f)$ is simple by the aforementioned Theorem 23 of [Smi17]. Clearly $\mathrm{Homeo}(\mathcal{B})^{\prime}$ is not trivial and it is a normal subgroup of $\mathrm{Ker}(f)$, so we can finally conclude that the index-$2$ simple subgroup $\mathrm{Ker}(f)$ of $\mathrm{Homeo}(\mathcal{B})$ is none other than $\mathrm{Homeo}(\mathcal{B})^{\prime}$. ∎

### v2

@@ Introduction
Thanks to this uniqueness, we denote by $\mathcal{A}$ the airplane and $\mathcal{R}_{n}$ the $n$-regular rabbit. The key tool to prove A is to associate another topological space that encodes the arrangements of circles: the dendrite or tree of circles. For the airplane, this dendrite is the universal Ważewski dendrite $D_{\infty}$. Let us recall that a dendrite is a Peano continuum where any two points are joined by a unique arc and can be thought as a compact metric tree. The universal Ważewski dendrite is the unique dendrite such that branch points have infinite order and are dense (it is universal in the sense that it includes homeomorphically any other dendrite, see [Waż23]). For the regular $n$-rabbit, the tree of circles is the $(n,\infty)$-biregular tree $T_{n,\infty}$, i.e., any edge joins a vertex of degree $n$ and one of infinite degree.
@@
In [BM00], Burger and Mozes introduced the concept of a universal group $\mathcal{U}(\Gamma)$ for automorphisms of the $n$-regular tree $T_{n}$, given a permutation group $\Gamma\leq\mathrm{Sym}([n])$ (where $n]=\{1,\dots,n\}$ if $n$ is finite and $\infty]=\mathbf{N}$). The universal group $\mathcal{U}(\Gamma)$ is characterized by the property that the local action, i.e., the action induced on the neighborhood of each vertex, lies in $\Gamma$. This construction was later generalized by Smith in [Smi17] for biregular trees $T_{m,n}$, where $m,n$ are countable and possibly infinite. In this case, two permutation groups $\Gamma_{m}\leq\mathrm{Sym}([m])$ and $\Gamma_{n}\leq\mathrm{Sym}([n])$ are used to construct the universal group $\mathcal{U}(\Gamma_{m},\Gamma_{n})$.
@@
Figure 5: The 44-rabbit lamination and the airplane lamination.
@@
The group $\mathrm{Homeo}^{+}(\mathcal{A})$ and the groups $\mathrm{Homeo}^{+}(\mathcal{R}_{n})$ are the subgroups of $\mathrm{Homeo}^{+}(\mathbf{S}^{1})$ that preserve the associated laminations, which are closed in $\mathrm{Homeo}^{+}(\mathbf{S}^{1})$.
Moreover, we prove in Section B.3 that the groups $\mathrm{Homeo}^{+}(\mathcal{A})$ and $\mathrm{Homeo}^{+}(\mathcal{R}_{n})$ correspond to the groups of homeomorphisms of the Julia sets that can be extended to homeomorphisms of the entire complex plane.
Consequences of the identification of the groups
@@
    $\mathrm{Homeo}^{+}(\mathcal{R}_{n})\hookrightarrow\mathrm{Homeo}^{+}(\mathcal{R}_{m})\hookrightarrow\mathrm{Homeo}^{+}(\mathcal{B}).$
In particular, the first point implies that
[…]
@@
The topological groups $\mathrm{Homeo}(\mathcal{R}_{n})$ and $\mathrm{Homeo}^{+}(\mathcal{R}_{n})$ are locally Roelcke precompact, in particular locally bounded, and quasi-isometric to the infinite-order tree $T_{\infty}$. On the other hand, the topological groups $\mathrm{Homeo}(\mathcal{A})$ and $\mathrm{Homeo}^{+}(\mathcal{A})$ are Roelcke precompact and thus coarsely bounded.
@@
In the case of $\mathrm{Homeo}(\mathcal{A})$, this group can be identified with a kaleidoscopic group where the local action is oligomorphic. As a result, the identification of the universal minimal flow of $\mathrm{Homeo}(\mathcal{A})$ follows directly from [BT23, Theorem 1.1]. For $\mathrm{Homeo}^{+}(\mathcal{A})$, however, the universal minimal flow has a particularly explicit description. Following [GM18], for a subset $A\subseteq\mathbf{S}^{1}$, let $\mathrm{Split}(\mathbf{S}^{1},A)$ denote the compact space obtained from $\mathbf{S}^{1}$ (with its continuous cyclic order) by replacing each point $a\in A$ with a pair of points $a^{-},a^{+}$, where $a^{+}$ is the cyclic successor of $a^{-}$. Let $P$ be the countable set that corresponds to the preimage of the set of cut points of $\mathcal{A}$ under the Carathéodory loop.
@@
gives an example where the universal minimal flow is cyclically ordered while some other minimal flows are not, since $\mathcal{A}$ and $D_{\infty}$ have no invariant cyclic order under $\mathrm{Homeo}^{+}(\mathbf{S}^{1})$ (Proposition 6.6). This answers a question of Glasner and Megrelishvili, [GM18, Question 4.9].
@@
The pictures of Julia sets displayed in this article were produced using Python code developed by the first author, which is available at the following link: https://gitlab.imo.universite-paris-saclay.fr/bruno.duchesne/julia-sets-in-python
@@ Acknowledgements
We are grateful to Matthieu Joseph, Eli Glasner and Michael Megrelishvili and Yuri Neretin for their comments on a preliminary version of this text.
We are thankful to the anonymous referee for their comments, which led to corrections, improvements and a significant reduction of typos.
@@ 1.1 Topological characterization of rabbits
Without condition 3, A would include Julia sets such as that for the map $z\mapsto z^{2}+c$ where $c\approx-1.311$, which is depicted in Figure 6.
@@
  • a regular point if it belongs to a circle and is not a cut point (equivalently thanks to Lemma 1.3 right below, if it belongs to a unique circle).
@@
Let $I=[a,b]$ be an arc in a rabbit $\mathcal{R}$ and let $p\in(a,b)\coloneqq[a,b]\setminus\{a,b\}$. If $p$ is a cut point, then it belongs to multiple circles by Corollary 1.5, so there is nothing to prove. If $p$ is not a cut point, then $\mathcal{R}\setminus\{p\}$ is connected and thus path-connected because $\mathcal{R}$ is a Peano continuum. It follows that there exists an arc $I^{*}$ that joins $a$ and $b$ while avoiding $p$. Then $I\cup I^{*}$ must include some circle that contains $p$. ∎
@@ 1.3 Legal coloring of the tree of circles
Let $\mathcal{R}$ be an $n$-regular rabbit. Let $\Omega$ be a countable set together with a dense cyclic order $O_{\Omega}$ and, if $n$ is finite, let $n]=\{1,\dots,n\}$ be equipped with its standard cyclic order. For each circle $C$ of $\mathcal{R}$, equip $\mathrm{out}(C)$ with a countable dense cyclic order on the cut points of $C$ as described in Remark 1.19 and, if $n$ is finite, for each cut point $p$ of $\mathcal{R}$, equip $\mathrm{out}(p)$ with a cyclic order. Then there is a legal coloring with colors $([n],\Omega)$ whose restriction to $\mathrm{out}(v)$ is an isomorphism of cyclic orders for each vertex $v$ (only for those that correspond to circles, if $n$ is infinite). In particular, by Lemma A.5 it is an isomorphism of the induced separation relations (Definition A.2).
@@
Suppose that $p$ is a cut point that does not belong to $T$ and is adjacent in $T_{\mathcal{R}}$ to some circle $C$ of $T$. We need to show that we can add $p$ to $T$ along with all circles that are adjacent to it and extend the coloring in such a way that the above conditions are satisfied. This will eventually color every vertex of $T_{\mathcal{R}}$. Given $p$ and $C$ as above, define $k(p,C)$ as $k(q,C)$ for any other cut point $q$ in $T$ that belongs to $C$ (in the first step, when $V(T)=\{C\}$, this choice is instead arbitrary). Then complete k(p,)k(p,-) to a bijection such that $k(p,C^{\prime})$ satisfies condition 2. Next, for each circle $C^{\prime}\neq C$ that is adjacent to $p$ define $k(C^{\prime},p)=k(C,p)$, so that condition 3 holds at the vertex $p$. For each $C^{\prime}$, color $\mathrm{out}(C^{\prime})$ in such a way that condition 1 holds. The map $k$ extended in this way satisfies all conditions, so we are done. ∎
@@ 1.4 Universal groups of biregular trees
$k|_{\mathrm{out}(g(v))}\circ g|_{\mathrm{out}(v)}\circ k|_{\mathrm{out}(v)}^{-1}\in\begin{cases}N&\text{if $v\in A$},\\ M&\text{if $v\in B$},\end{cases}$ (1)
@@
$\sigma_{g}(v)\coloneqq k|_{\mathrm{out}(g(v))}\circ g|_{\mathrm{out}(v)}\circ k|_{\mathrm{out}(v)}^{-1}.$
@@ 1.5 Uniqueness of the regular rabbits
The following proposition shows that there is a unique $n$-regular rabbit for each $n$ (up to homeomorphisms). Given a rabbit $\mathcal{R}$ and its tree of circles $T$, below we denote by $\pi\colon\mathcal{R}\to\overline{T}$ the map that sends each end point to the corresponding point of $\partial T$, each cut point to itself (as a vertex of $T$) and each regular point to the unique circle to which it belongs (as a vertex of $T$).
@@
Let $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ be two $n$-regular rabbits and let $T_{1}$ and $T_{2}$ be their trees of circles. Fix two legal colorings satisfying Lemma 1.21 and let $\varphi\colon T_{1}\to T_{2}$ be a color-preserving automorphism (the local action at any vertex is the identity). Note that the maps $\pi_{1}$ and $\pi_{2}$ are injective on the sets of cut points of $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$. This defines a bijection $\psi=\pi^{-1}_{2}\circ\varphi\circ\pi_{1}$ between the cut points of $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$. Since $\varphi$ is a tree isomorphism and the edge adjacency in $T_{1}$ and $T_{2}$ is given by the inclusion of cut points to circles, the map $\psi$ restricts to bijections between the cut points of each circle. Moreover, since $\varphi$ is color-preserving, the circular order of cut points on each circle is preserved, and since the sets of cut points are dense in $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ (by condition 1 of A together with Lemma 1.21), $\psi$ extends uniquely to a continuous map between a circle of $\mathcal{R}_{1}$ and its image. Doing this for all circles of $\mathcal{R}_{1}$ extends $\psi$ to a bijection between $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ such that $\varphi\circ\pi_{1}=\pi_{2}\circ\psi$. It now remains to show that $\psi$ is a homeomorphism.
@@ 1.6 Homeomorphism groups of regular rabbits
Since both are Polish groups, it suffices to prove that the abstract group isomorphism $\Pi\colon\mathrm{Homeo}(\mathcal{R})\to\mathcal{U}(\mathrm{Sym}([n]),\mathrm{Aut}(S))$ is continuous (for example because of [Gao09, Corollary 2.3.4]). Suppose that $g_{m}$ is a sequence in $\mathrm{Homeo}(\mathcal{R})$ that converges to the identity and consider an arbitrary circle $C_{0}$ of the rabbit $\mathcal{R}$. We will show that, for large enough $m\in\mathbf{N}$, the homeomorphism $g_{m}$ fixes $C_{0}$. Let $C_{1}$, $C_{2}$ and $C_{3}$ be circles of $\mathcal{R}$ that lie in three distinct connected components $U_{1}$, $U_{2}$ and $U_{3}$ of $\mathcal{R}\setminus C_{0}$. Then, for $m$ large enough, say $m\geq N$, the circle $g_{m}(C_{i})$ is included in $U_{i}$ for each $i=1,2,3$. In the tree of circles, consider the unique paths between $C_{1}$ and $C_{2}$, between $C_{1}$ and $C_{3}$ and between $C_{2}$ and $C_{3}$: by Remark 1.13, they intersect precisely at $C_{0}$, so the vertex $C_{0}$ is fixed by $\Pi(g_{m})$ for all $m\geq N$. This argument shows that each vertex is fixed by $\Pi(g_{m})$ for large enough $m$, so the sequence $\Pi(g_{m})$ converges to the identity of $\mathcal{U}(\mathrm{Sym}([n]),\mathrm{Aut}(S))$ and thus $\Pi$ is continuous. ∎
@@ 1.7 Consequences of the identification with universal groups
Consider now the case $n=2$, which concerns the basilica $\mathcal{R}_{2}=\mathcal{B}$. Neretin noted that $\mathrm{Homeo}(\mathcal{B})$ is the universal Burger-Mozes group $\mathcal{U}(\mathrm{Aut}(S))$ on a regular tree [Ner24], which has an index-$2$ simple subgroup generated by the edge stabilizers by [BM00, Proposition 3.2.1]. The regular tree in question is a simplified tree of circles whose vertices are the circles of the basilica and whose edges connect adjacent circles. Hence, the group generated by the pointwise stabilizers of edges corresponds to the group that preserves the alternating $2$-coloring of the circles depicted in Figure 8. This is the kernel of the group homomorphism $f\colon\mathrm{Homeo}(\mathcal{B})\to\mathrm{Sym}([2])=\{0,1\}$ that has value 0 or $1$ depending on whether the element preserves the coloring or not, which implies that $\mathrm{Homeo}(\mathcal{B})^{\prime}\leq\mathrm{Ker}(f)$. Moreover, note that $\mathrm{Ker}(f)=\mathcal{U}(1,\mathrm{Aut}(S))$, so $\mathrm{Ker}(f)$ is simple by the aforementioned Theorem 23 of [Smi17]. Clearly $\mathrm{Homeo}(\mathcal{B})^{\prime}$ is not trivial and it is a normal subgroup of $\mathrm{Ker}(f)$, so we can finally conclude that the index-$2$ simple subgroup $\mathrm{Ker}(f)$ of $\mathrm{Homeo}(\mathcal{B})$ is none other than $\mathrm{Homeo}(\mathcal{B})^{\prime}$. ∎